### v1

# Theory of In-Plane-Magnetic-Field-Dependent Excitonic Spectra in Atomically Thin Semiconductors

Abstract
The linear absorption spectrum of excitons in TMDC monolayers under the influence of an in-plane magnetic field is theoretically studied. We demonstrate that in-plane magnetic fields induce a hybridization between spin-bright and spin-dark exciton transitions, resulting in a brightening of spin-dark excitons. We analytically investigate spectral features including resonance energy shifts, broadening and amplitudes ratios. In particular, for a MoSe2 monolayer with radiatively-limited linewidth, we find a complex interplay of dark-bright splitting and linewidth difference of both involved spin-bright and spin-dark excitons.
TMDC Monolayer, Linear Absorption Spectrum, In-Plane Magnetic Field, Spin Hybridization, Non-Hermitian Quantum Mechanics
preprint: APS/123-QED
I Introduction
Transition metal dichalcogenides (TMDC) monolayers are atomically thin semiconductors of the $MX_{2}$ type, where $M=\{\text{Mo},\text{W}\}$ and $X=\{\text{S},\text{Se}\}$. The screening of the Coulomb interaction compared to their bulk forms is drastically reduced, which results in large excitonic binding energies of a few hundreds of meV. As a consequence, excitons dominate the linear optical response of such materials in the visible region [1, 2, 3].
Figure 1: (a) Spin-bright (light-gray ellipse) and spin-dark (dark-gray ellipse) excitonic transitions of the A-exciton considered in our model at the $K$-valleys. (b) Corresponding spin-bright and spin-dark transitions of the B-exciton. (c): An optical field $E_{0}^{\sigma}$ induces a polarization $P^{\sigma}$ in a TMDC monolayer in the presence of an in-plane magnetic field $B_{\parallel}$.
Due to strong spin-orbit-interaction-induced band splitting [4, 5], the valley and spin of the electrons in TMDC monolayers provide two degrees of freedom, cf. Fig. 1(a)–(b), which can be optically addressed, giving rise to possible applications in valleytronics [6] and spintronics [7]. As depicted in Fig. 1(a)–(b), excitons are called spin-bright if electron and hole assume a parallel spin-configuration and consequentially own a non-vanishing optical transition dipole moment. On the contrary, excitons with anti-parallel spin-configuration are called spin-dark, due to a vanishing optical transition dipole moment, and thus can not be seen in the absorption spectrum.
The manipulation of excitonic spin properties in atomically thin semiconductors magnetic fields is a fascinating and ongoing subject of research: Out-of-plane magnetic fields induce Zeeman- and diamagnetic shifts [8, 9, 10, 11] and excitonic Landau levels at high field-strengths [12, 13, 14]. In-plane magnetic fields induce spin coupling of spin-bright and spin-dark excitonic transitions via the spin angular momentum, cf. Fig. 1(a)–(b), which leads to the optical brightening of former spin-forbidden $s$-orbital states in photoluminescence (PL) or absorption [15, 16, 17, 18, 19, 20, 21, 22].
In a similar way, in-plane electric fields yield substantial Stark shifts of optically bright $s$-orbital states [23, 24, 25] and cause an angular-momentum mixing [26] in absorption spectra. On the other hand, out-of-plane electric fields cause a weak quantum-confined Stark effect [27] and spin coupling via Rashba spin-orbit interaction [28], which leads to the optical brightening of mixed spin-forbidden $s$- and $p$-orbital states [29, 30].
In this paper, we employ a thorough analytical approach to the effects of in-plane magnetic fields on the energetically lowest excitons in atomically thin semiconductors. We present full analytical expressions for the optical absorption, hybridized energies, linewidths and amplitudes including dissipative processes such as radiative decay or phonon-assisted decay. The analytical formulas can be easily used to analyze experimental spectra. We distinguish between two classes of materials: spin-dark materials such as MoS2, WS2 and WSe2 with lowest-lying excitonic states of unequal spin, which usually show an energetic separation between the spin-dark and spin-bright state exceeding 10 meV [20, 31, 32], and spin-bright materials such as MoSe2, which exhibit lowest-lying excitonic states of equal spin and a dark-bright splitting close to 1 meV [19, 20]. The paper is organized as follows: First, we develop the equations of motion for the spin-bright and spin-dark excitonic dipole (Sec. II) and their solutions (Sec. III) under the influence of an in-plane magnetic field, then, we calculate the absorption spectra of two example TMDC monolayers of the two different classes, MoSe2 and MoS2 (Sec. IV), and, at last, we analytically analyze the occurring magnetic-field-dependent spectral features (Sec. IV.1Sec. IV.3).
II Equations of Motion
To construct the absorption $\alpha(\omega)$ within the coupled Maxwell-Bloch formalism [33], we need the macroscopic polarization $P^{\,\sigma}$ induced by an external optical field $E_{0}^{\,\sigma}$ with polarization $\sigma=\sigma_{\pm}$, cf. Fig. 1(c). Since we consider an optical field, which strikes the sample perpendicularly, cf. Fig. 1(c), we restrict ourselves to intravalley transitions with vanishing in-plane center-of-mass momentum. In rotating-wave approximation, $P^{\,\sigma}$ reads:
$P^{\,\sigma}(t)=\sum_{\xi}(\delta_{\sigma,\sigma_{+}}\delta_{\xi,K}+\delta_{\sigma,\sigma_{-}}\delta_{\xi,K^{\prime}})\\ \times\sum_{s,\nu,\textbf{q}}\varphi_{\nu,\textbf{q}}^{\xi,s,s}(d_{\xi,s}^{\,c,v})^{*}P^{\,\xi,s,s}_{\nu}(t).$ (1)
Here, $P^{\,\xi,s,\xi,s}_{\nu}$ is the excitonic transition [34]:
$P^{\,\xi,s_{1},s_{2}}_{\nu}=\sum_{\mathbf{q}}\varphi_{\nu,\textbf{q}}^{*,\xi,s_{1},s_{2}}\langle\hat{v}^{\dagger,\xi,s_{1}}_{\mathbf{q}}\hat{c}^{\xi,s_{2}}_{\mathbf{q}}\rangle,$ (2)
where $\mathbf{q}$ are the relative momenta of the corresponding electron-hole pair with valence (conduction) band creation (annihilation) operator $\hat{c}^{\dagger}$ ($\hat{v}$) at valley $\xi$ and spin $s$. $\nu$ is the excitonic quantum number and $\varphi_{\nu,\mathbf{q}}^{\xi,s_{1},s_{2}}$ is the excitonic wave function solving the Wannier equation and $d_{\xi,s}^{\,c,v}$ is the absolute value of the transition dipole moment [35].
The exciton dynamics is obtained via Heisenberg’s equations of motion:
$\partial_{t}P^{\,\xi,s_{1},s_{2}}_{\nu}=\frac{i}{\hbar}\langle\big[\hat{H},\ \hat{P}^{\,\xi,s_{1},s_{2}}_{\nu}\big]\rangle,$ (3)
in the regime of linear optics. The total Hamiltonian
$\hat{H}=\hat{H}_{\text{X-0}}+\hat{H}_{\text{X-light}}+\hat{H}_{B_{\parallel}},$ (4)
consists of the free excitonic Hamiltonian $\hat{H}_{\text{X-0}}$, the exciton-light interaction Hamiltonian in length gauge $\hat{H}_{\text{X-light}}$ and the contribution due to a spatially homogeneous in-plane magnetic field interacting with the spin angular momentum $\hat{H}_{B_{\parallel}}$:
$\hat{H}_{B_{\parallel}}=\sum_{\begin{subarray}{c}\nu_{1},\nu_{2},\\ \xi,s_{1},s_{2}\end{subarray}}\left(\mathcal{B}_{\nu_{1},\nu_{2}}^{e,\xi,s_{1},s_{2}}\hat{P}^{\dagger}\vphantom{P}_{\nu_{1}}{}^{\xi,s_{1},s_{2}}\hat{P}\vphantom{P}_{\nu_{2}}^{\xi,s_{1},\bar{s}_{2}}\right.\\ \left.-\mathcal{B}_{\nu_{1},\nu_{2}}^{h,\xi,s_{1},s_{2}}\hat{P}^{\dagger}\vphantom{P}_{\nu_{1}}{}^{\xi,s_{1},s_{2}}\hat{P}\vphantom{P}_{\nu_{2}}^{\xi,\bar{s}_{1},s_{2}}\right).$ (5)
Here, quadratic contributions are neglected as the effect is already small in the linear regime and, to the best of our knowledge, no diamagnetic shifts have been measured yet with an in-plane geometry for TMDC monolayers. The in-plane interaction strength is governed by the excitonic magnetic matrix elements:
$\begin{split}\mathcal{B}_{\nu_{1},\nu_{2}}^{e,\xi,s_{1},s_{2}}&=\frac{g_{e}}{2}\mu_{B}B_{\parallel}\sum_{\textbf{q}}\varphi_{\nu_{1},\mathbf{q}}^{*\,\xi,s_{1},s_{2}}\varphi_{\nu_{2},\textbf{q}}^{\xi,s_{1},\bar{s}_{2}},\\ \mathcal{B}_{\nu_{1},\nu_{2}}^{h,\xi,s_{1},s_{2}}&=\frac{g_{e}}{2}\mu_{B}B_{\parallel}\sum_{\mathbf{q}}\varphi_{\nu_{1},\mathbf{q}}^{*\,\xi,s_{1},s_{2}}\varphi_{\nu_{2},\mathbf{q}}^{\xi,\bar{s}_{1},s_{2}},\end{split}$ (6)
where $\mu_{B}=\frac{e\hbar}{2m_{0}}$ is the Bohr magneton and $B_{\parallel}$ is the magnetic field strength. The $g$-factor can approximately be described by the free-electron $g$-factor $g_{e}\approx 2$ [5, 19, 22]. Note that $\overline{s}$ corresponds to the opposite spin of $s$, i.e., if $s=\,\uparrow$, then $\overline{s}=\,\downarrow$ and vice versa, which is caused by the action of an in-plane magnetic field on the spin wave functions. Therefore, the magnetic Hamiltonian is non-diagonal with respect to the out-of-plane spin states of the electrons and holes of the corresponding excitons. The consequences of this situation for linear optical spectra are the main focus of this paper. The magnetic matrix elements in Eq. 6 determine the spin-coupling selection rules: Due to the spatial homogeneity of the in-plane magnetic field $B_{\parallel}$, a coupling of angle-independent $s$-orbital states with angle-dependent $p$-orbital states does not occur. Moreover, a coupling between the $1s$-exciton and higher-lying $ns$-excitons quickly fades as $n$ increases due to a decreasing overlap of the corresponding exciton wave functions. Note that the former is a characteristic property of the interaction of a spatially homogeneous in-plane magnetic field with the spin angular momentum, as excitonic spin flips via spin-orbit interaction within spatially homogeneous [29, 36] or spatially inhomogeneous [37] out-of-plane electric fields, i.e., Rashba interaction [28], behave differently .
By using bosonic excitonic commutator relations [34] valid in the limit of linear optics, we obtain the following equations of motion for the excitonic transitions:
$\begin{split}\partial_{t}P_{\nu_{1}}^{\,\xi,s_{1},s_{2}}(t)=&\,-\Big(\frac{i}{\hbar}\varepsilon_{x,\nu_{1}}^{\xi,s_{1},s_{2}}+\gamma_{\nu_{1}}^{\xi,s_{1},s_{2}}\Big)P_{\nu_{1}}^{\,\xi,s_{1},s_{2}}(t)\\ &+i\Omega_{\nu}^{\xi,s_{1},s_{2}}(t)\\ &-\frac{i}{\hbar}\sum_{\nu_{2}}\mathcal{B}_{\nu_{1},\nu_{2}}^{e,\xi,s_{1},s_{2}}P_{\nu_{2}}^{\,\xi,s_{1},\bar{s}_{2}}(t)\\ &+\frac{i}{\hbar}\sum_{\nu_{2}}\mathcal{B}_{\nu_{1},\nu_{2}}^{h,\xi,s_{1},s_{2}}P_{\nu_{2}}^{\,\xi,\bar{s}_{1},s_{2}}(t).\end{split}$ (7)
The first line of Eq. 7 denotes the free excitonic contribution with excitonic energy $\varepsilon_{x,\nu_{1}}^{\xi,s_{1},s_{2}}$ and total dephasing:
$\gamma_{\nu}^{\xi,s_{1},s_{2}}=\gamma_{\text{nrad},\nu}^{\xi,s_{1},s_{2}}+\gamma_{\text{rad},\nu}^{\xi,s_{1},s_{2}},$ (8)
which consists of a non-radiative dephasing $\gamma_{\text{nrad}}^{\xi,s_{1},s_{2}}$ with a contribution due to exciton-phonon interaction, which can be calculated [38], or due to disorder [39] or strain [40], which are introduced phenomenologically, and a radiative contribution $\gamma_{\text{rad},\nu}^{\xi,s_{1},s_{2}}=\frac{\varepsilon_{x,\nu}^{\xi,s_{1},s_{2}}}{\hbar^{2}2\epsilon_{0}c_{0}n_{\text{refr}}}|d_{\xi,s_{1}}^{\,c,v}|^{2}\delta_{s_{1},s_{2}}$ due to reradiation [33] with vacuum permittivity $\epsilon_{0}$, speed of light $c_{0}$ and mean refractive index $n_{\text{refr}}=\frac{1}{2}(\sqrt{\epsilon_{1}}+\sqrt{\epsilon_{2}})$ of the homogeneous media surrounding the TMDC monolayer from both sides. Note, that we assume $\epsilon_{1}=\epsilon_{2}$ throughout this work. The second line describes the coupling to an optical field via the excitonic Rabi frequency:
$\begin{split}\Omega_{\nu}^{\xi,s_{1},s_{2}}(t)=&\,\frac{1}{\hbar}\sum_{\textbf{q}}\varphi_{\nu,\textbf{q}}^{*\,\xi,s_{1},s_{2}}d_{\xi,s_{1}}^{\,c,v}\\ &\times\sum_{\sigma}(\delta_{\xi,K}\delta_{\sigma,\sigma_{+}}+\delta_{\xi,K^{\prime}}\delta_{\sigma,\sigma_{-}})E_{0}^{\sigma}(t),\end{split}$ (9)
which depends on the dipole matrix element $d^{\,c,v}_{\xi,s_{1}}$, the incident optical field $E_{0}^{\sigma}(t)$ with polarization $\sigma$ and the Coulomb enhancement $\sum_{\textbf{q}}\varphi_{\nu,\textbf{q}}^{*\,\xi,s_{1},s_{2}}$. The third and fourth line of Eq. 7 describes the coupling of an excitonic transition with quantum number $\nu_{1}$ to all other transitions with quantum number $\nu_{2}$ via electron spin flips (third line) and hole spin flips (second line) due to the effect of the in-plane magnetic field $B_{\parallel}$.
III Hybridized Excitonic States
Without loss of generality, in the following, we restrict ourselves to right-handed polarization $\sigma=\sigma_{+}$ that can optically excite electron-hole pairs at the $K$-valley only. Since we work in the regime of linear optics, our description is equally valid for left-handed or linear optical excitation and we can neglect intervalley coupling mechanisms, which are mostly relevant in nonlinear dynamics [41, 42, 43, 44, 45, 46, 47] or states with non-vanishing center-of-mass momenta Q [48, 49]. Moreover, we restrict the description to the lowest-lying $1s$-excitons of the A-series, where we neglect A/B-coupling, since it is strongly suppressed due to the large valence band splitting of several hundreds of meV [4]. The remaining two equations of motion from Eq. 7 can be solved analytically in frequency space by first eliminating the spin-dark ($d$) excitonic transition in the emerging analytical expression of the spin-bright ($b$) excitonic transition. The spin-bright transition $P^{b}(\omega)=P_{1s}^{\xi,s,s}$, which couples directly to the Maxwell field via Eq. 1 and contains – due to spin mixing – also spectral signatures of the dark transition, then reads:
$P^{b}(\omega)$ $=\frac{\hbar\Omega^{b}(\omega)\big(\hbar\omega_{x}^{d}-\hbar\omega-i\hbar\gamma^{d}\big)}{\big(\hbar\omega_{x}^{b}-\hbar\omega-i\hbar\gamma^{b}\big)\big(\hbar\omega_{x}^{d}-\hbar\omega-i\hbar\gamma^{d}\big)-\mathcal{B}^{2}}$ (10)
Here, $b=\{K,\uparrow,\uparrow,1s\}$ denotes spin-bright and $d=\{K,\downarrow,\uparrow,1s\}$ denotes spin-dark with respect to the initial excitonic configuration at $B_{\parallel}=0$, cf. Fig. 1(a)–(b) and we define $\mathcal{B}=|\mathcal{B}_{\nu}^{e,K,\uparrow,\uparrow}|$, cf. Eq. 6. Subsequently, we can cast the equation in a more intuitive form by performing a partial fraction decomposition:
$P^{b}(\omega)=\hbar\Omega^{b}(\omega)\sum_{\mathcal{S}}\frac{P^{\mathcal{S}}_{B_{\parallel}}}{\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega-i\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}}.$ (11)
As a consequence of the magnetic-field induced spin coupling, the spin-bright transition in Eq. 11 is now a superposition of two excitonic resonances, the hybridized states labeled by the spin-diagonal quantum number $\mathcal{S}\in\{1,-1\}$ centered around their respective hybridized energies:
$\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}=\frac{1}{2}(\hbar\omega_{x}^{b}+\hbar\omega_{x}^{d})\\ +\mathcal{S}\frac{1}{2\sqrt{2}}\sqrt{\mathcal{C}(\mathcal{B};\Delta;\kappa)-\kappa^{2}+\Delta^{2}+4\mathcal{B}^{2}},$ (12)
with hybridized linewidths:
$\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}=\frac{1}{2}(\hbar\gamma^{d}+\hbar\gamma^{b})\\ +\mathcal{S}\frac{\text{sgn}(\Phi)}{2\sqrt{2}}\sqrt{\mathcal{C}(\mathcal{B};\Delta;\kappa)+\kappa^{2}-\Delta^{2}-4\mathcal{B}^{2}},$ (13)
We define:
$\mathcal{C}(\mathcal{B};\Delta;\kappa)=\\ \sqrt{8\mathcal{B}^{2}\Big(-\kappa^{2}+\Delta^{2}+2\mathcal{B}^{2}\Big)+\Big(\kappa^{2}+\Delta^{2}\Big)^{2}},$ (14)
with dark-bright splitting $\Delta$:
$\Delta=|\hbar\omega_{x}^{b}-\hbar\omega_{x}^{d}|,$ (15)
and linewidth difference $\kappa$:
$\kappa=|\hbar\gamma^{b}-\hbar\gamma^{d}|.$ (16)
The quantity:
$\Phi=(\omega^{b}_{x}-\omega^{d}_{x})(\gamma^{b}-\gamma^{d}),$ (17)
determines the sign in Eq. 13 and ensures correct convergence behavior in the limit $B_{\parallel}\rightarrow 0$.
The amplitudes of both occurring resonances in Eq. 11 are scaled with the complex mixing coefficients $P^{\mathcal{S}}_{B_{\parallel}}$:
$P^{\mathcal{S}}_{B_{\parallel}}=\frac{\omega_{x}^{d}-i\gamma^{d}-\omega_{x,B_{\parallel}}^{\mathcal{S}}+i\gamma_{B_{\parallel}}^{\mathcal{S}}}{\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}}-i\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}}-\omega_{x,B_{\parallel}}^{\mathcal{S}}+i\gamma_{B_{\parallel}}^{\mathcal{S}}},$ (18)
which obey the following identity
$P^{\mathcal{S}}_{B_{\parallel}}+P^{\overline{\mathcal{S}}}_{B_{\parallel}}=1.$ (19)
We note that due to the explicit inclusion of different linewidths of the involved spin-bright and spin-dark excitons, which directly impact the spin hybridization, our excitonic description resembles a non-Hermitian treatment to the optical response of atomically thin semiconductors [50]. These non-Hermiticy-resembling effects are most pronounced, if energy splitting, linewidth and linewidth differences between the involved resonances are all of similar magnitude. Such a regime can occur in, e.g., disorder-/defect-free h-BN-encapsulated MoSe2 monolayers at cryogenic temperatures with a dark-bright energy splitting $\Delta$ of 1.5 meV and an almost radiatively-limited linewidth $\hbar\gamma$ of the spin-bright exciton of approximately 111.51.5\,meV.
The mixing coefficients from Eq. 18 are plotted as an example for small non-radiative linewidth in Fig. 2. Here, the real part redistributes the oscillator strength with increasing magnetic field between both initial spin-bright and spin-dark excitonic transitions and therefore displays the degree of spin hybridization: At zero magnetic field, no redistribution takes place and the mixing coefficient corresponding to the spin-dark excitonic transition is zero. At an increasing magnetic field, the mixing coefficient of the initial spin-bright transition decreases and the mixing coefficient of the initial spin-dark transition increases, until both mixing coefficients converge to
$\lim_{B_{\parallel}\mapsto\infty}P^{\mathcal{S}}_{B_{\parallel}}=\lim_{B_{\parallel}\mapsto\infty}P^{\overline{\mathcal{S}}}_{B_{\parallel}}=0.5,$ (20)
for very large magnetic field strengths. However, this regime clearly exceeds currently experimentally available field strengths on the order of 100 T [12]. In Eq. 18 and Fig. 2, additional imaginary components appear because we do not neglect dissipation throughout our whole analysis.
Figure 2: Real- and imaginary part of mixing coefficients $P_{B_{\parallel}}^{\mathcal{S}}$ for $\nu=1s$ under influence of the in-plane magnetic field for MoSe2 ($\hbar\gamma_{\text{nrad}}=0.5$ meV).
Note that because of the varying energy ordering of spin-bright and spin-dark excitons in the respective sample of interest, the assignment of the $\mathcal{S}$-states to specific spin-bright or spin-dark states can change.
IV Linear Absorption Spectrum
The linear absorption spectrum $\alpha(\omega)$ can be expressed as [51]:
$\alpha(\omega)=1-T(\omega)-R(\omega),$ (21)
where $T(\omega)$ and $R(\omega)$ are the transmission and reflection coefficients, respectively:
$T(\omega)$ $=\bigg|\frac{\textbf{E}_{T}(\omega)}{\textbf{E}_{0}(\omega)}\bigg|^{2}=\Bigg|\frac{\tilde{E}_{0}^{\sigma_{+}}(\omega)+\frac{i\varepsilon^{b}_{x}}{2\hbar\epsilon_{0}c_{0}n_{\text{refr}}}\tilde{P}^{\sigma_{+}}(\omega)}{\tilde{E}^{\sigma_{+}}_{0}(\omega)}\Bigg|^{2},$ (22)
$R(\omega)$ $=\bigg|\frac{\textbf{E}_{R}(\omega)}{\textbf{E}_{0}(\omega)}\bigg|^{2}=\Bigg|\frac{\frac{i\varepsilon^{b}_{x}}{2\hbar\epsilon_{0}c_{0}n_{\text{refr}}}\tilde{P}^{\sigma_{+}}(\omega)}{\tilde{E}^{\sigma_{+}}_{0}(\omega)}\Bigg|^{2},$ (23)
that are computed with the macroscopic polarization $P^{\sigma}(t)$ from Eq. 1. By performing a partial fraction decomposition again and, with the help of Eq. 11, the absorption can be expressed explicitly as:
$\begin{split}\alpha(\omega)=\sum_{\mathcal{S}}&\frac{\mathcal{F}^{\mathcal{S}}-\mathcal{G}^{\mathcal{S}}\hbar\omega}{(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega)^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}},\end{split}$ (24)
which is a superposition of the two peak signals modulated by the amplitudes $\mathcal{F}_{\mathcal{S}},\mathcal{G}_{\mathcal{S}}$, cf. App. A. The amplitudes $\mathcal{F}_{\mathcal{S}}$ correspond to symmetric Lorentzian signals, which are modified by the amplitudes $\mathcal{G}_{\mathcal{S}}$ scaling linearly with the energy $\hbar\omega$. The latter are interference effects, which occur as the two hybridized excitonic frequencies come energetically close, causing an asymmetric deformation of the spectrum.
We depict in Fig. 3 the linear absorption spectra Eq. 24 of monolayers MoSe2 (left) and MoS2 (right) encapsulated in h-BN for three in-plane magnetic field-strengths around the spin-bright 1s-orbital state. From top to bottom, we vary the non-radiative linewidth, which describes the regime of a radiatively-limited linewidth up to a regime, where phonons gain a dominating influence (100 K) [38].
Figure 3: Linear absorption spectrum under influence of an in-plane magnetic field for $\nu=1s$ with different field-strengths for the material MoSe2 in (a)–(c) and MoS2 in (d)–(f) encapsulated in h-BN with $\sigma_{+}$-polarised light for increasing non-radiative linewidths $\gamma_{\text{nrad}}$.
To obtain Fig. 3, we adjusted the excitonic energies according to recent PL experiments with dark-bright splitting of 1.45 meV in the optically bright MoSe2 monolayer [19, 20] and dark-bright splitting of 14 meV in the optically dark MoS2 monolayer [20]. We calculated the Coulomb enhancement $\sum_{\textbf{q}}\varphi_{1s,\textbf{q}}^{*\,K,\uparrow,\uparrow}$ in Eq. 9 mediated by the excitonic wave functions as solutions of the corresponding Wannier equations and the transitions dipole moments with $\mathbf{k}\cdot\mathbf{p}$-parameters according to Refs. [35, 4]. All parameters are provided in Tab. 1 in the appendix.
In the absorption of a MoSe2 monolayer in Fig. 3 at increasing in-plane magnetic field strengths, the initially spin-dark resonance brightens, i.e., gains an oscillator strength. Simultaneously, both excitonic resonances shift in the opposite direction, i.e., an anti-crossing occurs. This has also been observed in PL [20, 52, 19, 53, 54, 21]. In the MoS2 monolayer, the same behavior occurs, but strongly suppressed due to the much larger dark-bright splitting of 14.5 meV compared to 1.45 meV in the MoSe2 monolayer. The brightening of initially spin-dark excitons is most pronounced in the radiatively-limited case of the linewidth, where a strong signal at the spin-dark resonance already appears at 15 T. At a slightly increased non-radiative linewidth of $\gamma_{\text{nrad}}=0.5$ meV, a clear spin-dark resonance appears at 30 T and a shoulder can be observed at 15 T in a MoSe2 monolayer, while, in a MoS2 monolayer, the initial spin-dark resonance is barely detectable even at 30 T. At comparably large non-radiative linewidths of $\gamma_{\text{nrad}}=5\,$meV, both resonances can no longer be separated and appear as one single, broadened peak.
Also, difficult to detect by eye, but present, is a redistribution of the linewidths in an increasing magnetic field, which is modelled by Eq. 13.
These three magnetic-field-dependent features – the peak position, the linewidth and the behaviour of the amplitude – are discussed in detail in the following.
IV.1 Hybridized Excitonic Energies
The hybridized excitonic energies $\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}$ are governed by Eq. 12 and plotted for the materials MoSe2 in Fig. 4(a) and MoS2 in Fig. 4(b) as a function of magnetic field $B_{\parallel}$ and linewidth difference $\kappa$ from Eq. 16.
(a) MoSe2
(b) MoS2
Figure 4: Hybridized excitonic energies $\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}$ from Eq. 12 for MoSe2 in Fig. 4(a) and MoS2 in Fig. 4(b) under influence of an in-plane magnetic field for increasing linewidth difference $\kappa$ from Eq. 16. The colourgradient of the curves show the values of the real parts of mixing coefficients from Eq. 18, i.e. the hybridization coefficients. Red (blue) represents a spin-up (spin-down) electron-state and yellow denotes the degree of spin-mixing.
As expected, the hybridized excitonic energies repel each other under influence of the in-plane magnetic field. In all cases, the initial growth as a function of $B_{\parallel}$ is quadratic. This can be seen analytically by performing a Taylor approximation at $B_{\parallel}=0$, yielding
$\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}\approx\frac{1}{2}(\hbar\omega_{x}^{b}+\hbar\omega_{x}^{d})+\mathcal{S}\Big(\frac{1}{2}\Delta\,+\mathcal{B}^{2}\frac{\Delta}{\kappa^{2}+\Delta^{2}}\Big).$ (25)
Here, for small $B_{\parallel}$-fields, the excitonic hybridized frequency grows quadratically in the field $B_{\parallel}$ and linear in the dark-bright splitting $\Delta$, renormalized by both the dark-bright splitting and the linewidth difference. The monolayer MoS2 in Fig. 4(b) does not leave this quadratic regime, because the excitonic magnetic matrix element still remains relatively small compared to the dark-bright splitting even up to 30 T.
On the contrary, for very high magnetic-field strengths and small dark-bright splittings, the growth becomes linear, which can be observed in the case for MoSe2 in Fig. 4(a). This regime can be analytically obtained from Eq. 12, if one assumes that the in-plane magnetic field dominates:
$\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}\approx\frac{1}{2}(\hbar\omega_{x}^{b}+\hbar\omega_{x}^{d})+\mathcal{S}\mathcal{B}.$ (26)
In the following, we discuss the impact of the linewidth on the level repulsion. The dependence on the linewidth difference is particularly pronounced for MoSe2 in Fig. 4(a), because it exhibits a small dark-bright splitting. As indicated by the color gradient, the spin-hybridization at a fixed magnetic field strength varies, if the linewidth difference is varied: If $\kappa$ increases, the degree of hybridization and hence the splitting decreases. Therefore, the dissipation of the system effectively suppresses spin-hybridization and can also be viewed as an effective attractive interaction.
To illustrate this situation, we distinguish two analytical limit cases. (i) In the case for dominating linewidth differences, we can eliminate the dark-bright-splitting-dependency by assuming that $\kappa\gg\Delta$ in the model Eq. 25, yielding:
$\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}\approx\frac{1}{2}(\hbar\omega_{x}^{b}+\hbar\omega_{x}^{d})+\mathcal{S}\Big(\frac{1}{2}\Delta\,+\mathcal{B}^{2}\frac{\Delta}{\kappa^{2}}\Big).$ (27)
In this case, the initial quadratic growth becomes exclusively inversely proportional to the square of $\kappa$. This holds true for MoSe2 in Fig. 4(a). (ii) On the other hand, when the dark-bright splitting $\Delta$ becomes the dominating parameter, which holds true for the majority of the TMDCs (MoS2, WS2, WSe2), we obtain an expression that does not depend on $\kappa$ anymore:
$\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}\approx\frac{1}{2}(\hbar\omega_{x}^{b}+\hbar\omega_{x}^{d})+\mathcal{S}\Big(\frac{1}{2}\Delta\,+\frac{\mathcal{B}^{2}}{\Delta}\Big),$ (28)
i.e., the influence of the linewidth difference vanishes. This regime is valid for MoS2 in Fig. 4(b), where the linewidth difference $\kappa$ only influences the shifts of energy marginally. At last, we emphasize the limit of $\kappa\approx 0$:
$\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}=\frac{1}{2}(\hbar\omega_{x}^{b}+\hbar\omega_{x}^{d})+\frac{\mathcal{S}}{2}\sqrt{\Delta^{2}+4\mathcal{B}^{2}}$ (29)
This case, for both materials in Figs. 4(a) and 4(b), leads to the strongest shifts, as the attractive part of the coupling induced by the dissipation is absent. Eq. 29 is the usual expression of level repulsion [55] and has been already discussed in other works [19, 18]. We note, that the magnetic-field-induced hybridization also manifests in time domain via quantum beats [56].
IV.2 Hybridized Excitonic Linewidths
The hybridized linewidths $\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}$ are governed by Eq. 13 and plotted for MoSe2 in Fig. 5. We note, that the linewidths discussed in this manuscript always correspond to a coherence decay via, e.g., radiative or pure dephasing and are not equal to a population decay due to radiative or nonradiative recombination [57].
Figure 5: Hybridized linewidth $\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}$, normalized w.r.t. summed linewidth $\hbar\gamma^{b}+\hbar\gamma^{d}$, in dependence of an in-plane magnetic field for MoSe2 with $\kappa$ = 0.71 meV and $\hbar\gamma_{\text{nrad}}=0.1$ meV. The colourgradient of the curves show the values of the real parts of mixing coefficients from Eq. 18, i.e. the hybridization coefficients. Red (blue) represents a spin-up (spin-down) electron-state and yellow denotes the degree of spin-mixing.
At increasing magnetic fields, it can be seen from Fig. 5 that the linewidths of the two hybridized states grow towards each other, which results in the fact – in contrast to the magnetic-field-dependent behavior of the hybridized excitonic energies, cf. Fig. 4 – that the hybridized linewidths approach their mean value at very high magnetic fields:
$\lim_{B_{\parallel}\mapsto\infty}\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}=\frac{1}{2}(\hbar\gamma^{b}+\hbar\gamma^{d}).$ (30)
In general, the resonance with the larger linewidth distributes a fraction to the resonance with the smaller linewidth. To demonstrate this in more detail, we depict the absolute difference of the hybridized linewidths in Fig. 6:
$|\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}-\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}}|=\frac{1}{\sqrt{2}}\sqrt{\mathcal{C}(\mathcal{B};\Delta;\kappa)+\kappa^{2}-\Delta^{2}-4\mathcal{B}^{2}}.$ (31)
For vanishing field-strength, it holds:
$|\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}-\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}}|\Big|_{B_{\parallel}=0}$ $=\kappa,$ (32)
cf. Fig. 6 at $B_{\parallel}=0\,$T. Therefore, if no linewidth difference is present, i.e. $\kappa=0$, there is no linewidth to redistribute and the impact of the magnetic field on the individual linewidths vanishes. Analytically, this follows from Eq. 13 directly via:
$\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}\Big|_{\kappa=0}=\hbar\gamma^{b}=\hbar\gamma^{d},$ (33)
Figure 6: Absolute difference of hybridized linewidths calculated with Eq. 31 for MoSe2 under influence of of an in-plane magnetic field for increasing linewidth differences $\kappa$
which is depicted in Fig. 6 for the $\kappa=0$ - case. Furthermore, for increased linewidth differences, Fig. 6 highlights that the redistribution of linewidth becomes stronger in terms of absolute values, yet needs stronger fields to converge to the mean value as the hybridization coefficients grows weaker as discussed in Sec. III. Similarly to Sec. IV.1, further limiting cases of Eq. 13 can be constructed, whose derivation we omit here.
As in Fig. 4, the hybridized linewidths in a MoSe2 monolayer, cf. Fig. 5, are much more sensitive to the magnetic field compared to MoS2, WS2 and WS2 monolayers due to its smaller dark-bright splitting.
IV.3 Absorption ratios of hybridized exciton peaks
By evaluating the absorption spectrum in Eq. 24 at the hybridized excitonic energies Eq. 12, we obtain the amplitude peak ratio of spin-dark and spin-bright resonances as:
$\begin{split}&\text{Peak~ratio}=\frac{\alpha(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})}{\alpha(\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})}.\end{split}$ (34)
In Fig. 7(a) and Fig. 7(b), we show the peak amplitude ratios from Eq. 34 for the materials MoS2 and MoSe2, respectively. We evaluate Eq. 34 in such a way, that the numerator (denominator) always corresponds to the spin-dark (spin-bright) resonance, and we choose a regime of quasi radiatively-limited linewidth, which corresponds to disorder-/defect-free h-BN-encapsulated samples at cryogenic temperatures.
(a) MoS2
(b) MoSe2
Figure 7: Amplitude peak ratios and spin-bright/spin-dark amplitudes, that are normalized w.r.t. $\mathcal{N}=\alpha(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})+\alpha(\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})$, for MoS2 (Fig. 7(a)) and MoSe2 (Fig. 7(b)) with non-radiative linewidth $\hbar\gamma_{\text{nrad}}=0.1$ meV for increasing in-plane magnetic field-strengths. The colourgradient of the curves show the values of the real parts of mixing coefficients from Eq. 18, i.e. the hybridization coefficients. Red (blue) represents a spin-up (spin-down) electron-state and yellow denotes the degree of spin-mixing.
In Fig. 7(a), the peak ratio of MoS2 (olive solid line) displays a monotonous increase within a magnetic field strength interval of up to 30 T. This behavior reflects the increased optical brightening of the initial spin-dark exciton, cf. also Fig. 4(b) and Fig. 2. We can derive an analytical limit case for Eq. 34 by utilizing the fact that the dark-bright splitting dominates over the linewidths. We evaluate numerator and denominator separately by using Eq. 28 and a similar limiting case for Eq. 13 and keeping only the terms that grow quadratically in the magnetic field strength and find:
$\begin{split}&\text{Peak~ratio}\,\Bigg|_{\text{MoS}_{2}}\approx\\ &\frac{(\hbar\gamma^{b})^{2}}{(\hbar\gamma^{d})^{2}}\frac{\mathcal{B}^{2}}{(\hbar\gamma^{b}-\hbar\gamma_{\text{rad}}^{b})\Delta^{2}+[\hbar\gamma^{d}+2\frac{\hbar\gamma^{b}}{\hbar\gamma^{d}}(\hbar\gamma^{b}-\hbar\gamma_{\text{rad}}^{b})]\mathcal{B}^{2}}.\end{split}$ (35)
This approximate expression, cf. silver dashed line in Fig. 7(a), describes the initial grow with respect to the magnetic-field strength, which reproduces the full expression, cf. olive solid line, very well. For magnetic fields above 30 T, however, LABEL:eq:_peak_ratio_MoS2 fails to reproduce the correct saturation behavior: It asymptotically approaches a ratio exceeding 50 $\%$, whereas the exact expression exhibits a maximum before converging to 50 $\%$ as illustrated for MoSe2 in Fig. 7(b). We further plot the quadratic approximation of LABEL:eq:_peak_ratio_MoS2, cf. silver dotted line in Fig. 7(a), with respect to the excitonic magnetic matrix element $\mathcal{B}^{2}$ around the origin, which reproduces the full expression (olive solid line) up to roughly 15 T. We note that the limit case in LABEL:eq:_peak_ratio_MoS2 can be used to model any material that exhibits a dominating dark-bright splitting over the linewidths, for example WS2 and WSe2 at cryogenic temperatures.
In contrast, within the field strengths considered here, the peak amplitude ratio for a MoSe2 monolayer, cf. olive solid line in Fig. 7(b), displays a strong non-monotonous behavior at increasing magnetic-field strengths. Initially, it increases, until a turning point at around 13 T is reached, from where it then slowly decreases. This non-monotonous behavior is a direct consequence of the fact that in a MoSe2 monolayer, dark-bright splitting, linewidths and linewidths differences are all of similar magnitude in the range of 1–2 meV. The strong initial growth of the peak ratio well beyond 100 $\%$ can be explained by the fact that, due to the absence of a radiative contribution, the linewidth of the spin-dark resonance is smaller compared to the spin-bright resonance. This in turn overemphasizes the oscillator strength it receives from hybridization (smaller linewidths lead to larger peak amplitudes at a constant area under the curve). For a closer analysis, we rewrite the peak ratio from Eq. 34 as:
$\begin{split}&\text{Peak~ratio}\Big|_{\text{MoSe}_{2}}=\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}g^{\mathcal{S}=1},\end{split}$ (36)
which is a product of a monotonously decreasing ratio of the hybridized linewidths $\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}$ and a monotonously increasing function $g^{\mathcal{S}}$, defined in Eq. 42, with respect to the magnetic field $B_{\parallel}$. Hence, the maximum of the amplitude ratio in Fig. 7(b) can be rationalized as that point, where the linewidth ratio $\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}$ reflecting the magnetic-field-induced linewidth redistribution starts to grow faster than the oscillator-strength redistribution encoded in $g^{\mathcal{S}}$. At magnetic fields exceeding the turning point, a decrease of the amplitude ratio occurs, until it is fully determined by the degree of the spin hybridization, as $\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}\overset{B_{\parallel}\rightarrow\infty}{\longrightarrow}1$ and $g^{\mathcal{S}}\overset{B_{\parallel}\rightarrow\infty}{\longrightarrow}0.5$. Consequently, if the linewidths are initially equal, the linewidth ratio becomes unity and the the peak ratio in Eq. 36 will be a purely monotonic increase with no extremum, as in the case for MoS2 in LABEL:eq:_peak_ratio_MoS2 (within the field strengths considered).
Note, that within our definition, cf. Eq. 34, the spin-dark amplitude and, thus, the ratio are not exactly zero for a vanishing in-plane magnetic field, cf. Fig. 7(b), because the spin-dark resonance lies in close vicinity to the spin-bright resonance.
We point out, that the qualitative behavior of the peak ratios resembles the behavior of the peak ratios extracted from PL experiments [20, 19]. However, while similar in their outer appearance, the underlying physics is different: In contrast to absorption measurements, which display the coherent response of the material, PL displays the emission of radiation. Hence, the PL amplitudes always scale with the incoherent excitonic occupations at the respective exciton resonance, which recombine within the light cone [57, 58, 52, 32], and the peak ratio can be rationalized as the ratio of the Boltzmann-distributed spin-bright and spin-dark occupations [19]. Therefore, future work can possibly shed more light on this issue, i.e., on the question, if the non-monotonous behavior can also be observed in actual absorption measurements.
Conclusion
We derived a fully analytical model for the absorption spectrum of the excitonic spin-dark and spin-bright ground states under spin hybridization of an applied in-plane magnetic field. We found, that an intricate interplay of hybridized excitonic energies, linewidths and amplitudes governs the linear optical response. In particular, we showed that a MoSe2 monolayer with quasi radiatively-limited linewidths gives rise to a non-monotonous amplitude behavior of spin-bright and spin-dark excitonic states in optical absorption spectra, similar to recent photoluminescence experiments. The derived equations Eqs. 24, 12 and 13, can be used to analyze magneto-optical experiments in the field of TMDC excitons.
Acknowledgements.
The authors acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG) through Project No. 420760124.
Appendix A Analytical Formulas
In this appendix, we show how we arrive at the absorption formula in Eq. 24 by introducing a Lorentzian of the spin-hybridized state $\mathcal{S}$ as
$L_{\mathcal{S}}(\omega)$ $=P^{\mathcal{S}}_{B_{\parallel}}\frac{i\hbar\gamma_{\text{rad}}^{b}}{\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega-i\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}}.$ (37)
With this definition, we can express the absorption spectrum from Eq. 21 as follows:
$\begin{split}\alpha(\omega)&=1-|L_{\mathcal{S}=1}(\omega)+L_{\mathcal{S}=-1}(\omega)|^{2}\\ &\qquad-|1+L_{\mathcal{S}=1}(\omega)+L_{\mathcal{S}=-1}(\omega)|^{2}\\ &\qquad-4\Re[L_{\mathcal{S}=1}(\omega)L_{\mathcal{S}=-1}^{*}(\omega)].\end{split}$ (38)
We plug in the definition from Eq. 37 and reorganize the expression using partial fraction decomposition, which leads to the absorption spectrum given in Eq. 24. The amplitudes are given by:
$\mathcal{F}^{\mathcal{S}}$ $=-2\Big[|(i\hbar\gamma_{\text{rad}}^{b})P^{\mathcal{S}}_{B_{\parallel}}|^{2}$ (39)
$\qquad+\Re[(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}+i\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})(P^{\mathcal{S}}_{B_{\parallel}}i\hbar\gamma_{\text{rad}}^{b}+2\mathcal{S}\Lambda)]\Big],$
$\mathcal{G}^{\mathcal{S}}$ $=-2\Re\Big[i\hbar\gamma_{\text{rad}}^{b}P^{\mathcal{S}}_{B_{\parallel}}+2\mathcal{S}\Lambda\Big],$ (40)
with:
$\Lambda=\frac{(\hbar\gamma_{\text{rad}}^{b})^{2}P^{\mathcal{S}=1}_{B_{\parallel}}P^{*\,\mathcal{S}=-1}_{B_{\parallel}}}{\hbar\omega_{x,B_{\parallel}}^{{\mathcal{S}=-1}}+i\hbar\gamma_{B_{\parallel}}^{{\mathcal{S}=-1}}-\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}=1}+i\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1}}.$ (41)
The function $g^{\mathcal{S}}$ from Eq. 36 reads:
$\begin{split}g^{\mathcal{S}}=&\frac{\big((\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}\big)}{\big((\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}}-\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}})^{2}\big)}\\ &\times\frac{(\mathcal{F}^{\mathcal{S}}-\mathcal{G}^{\mathcal{S}}\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})\big((\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}}-\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}})^{2}\big)+(\mathcal{F}^{\overline{\mathcal{S}}}-\mathcal{G}^{\overline{\mathcal{S}}}\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}}{(\mathcal{F}^{\mathcal{S}}-\mathcal{G}^{\mathcal{S}}\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})(\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}})^{2}+(\mathcal{F}^{\overline{\mathcal{S}}}-\mathcal{G}^{\overline{\mathcal{S}}}\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})\big((\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}\big)}.\end{split}$ (42)
Appendix B Parameters
Quantity MoSe2 MoS2
$1s$ spin-bright excitonic energy $\hbar\omega_{x,1s}^{b}$ 1.639 eV (4 K) [20] 1.931 eV (4 K) [20]
Dark-bright splitting $\hbar\omega_{x,1s}^{b}-\hbar\omega_{x,1s}^{d}$ 1.45-1.45 meV [19, 20] 14.5 meV [20, 31]
Effective spin-up electron mass $m_{e}^{K,\uparrow}$ [4] $0.5m_{0}$ $0.44m_{0}$
Effective spin-down electron mass $m_{e}^{K,\uparrow}$ [4] $0.58m_{0}$ $0.47m_{0}$
Effective spin-up hole mass $m_{h}^{K,\uparrow}$ [4] $0.6m_{0}$ $0.54m_{0}$
$\mathbf{k}\cdot\mathbf{p}$-parameter $\gamma_{\mathbf{k}\cdot\mathbf{p}}$ [4] $\frac{1}{2}(0.253+0.220)$ eV nm $\frac{1}{2}(0.276+0.222)$ eV nm
Transition dipole moment $d^{cv}=\frac{e\sqrt{2}\gamma_{\mathbf{k}\mathbf{p}}}{\tilde{E}_{\text{gap}}}$ 0.1689 $e$ nm 0.1536 $e$ nm
Monolayer width $d$ [59] 0.6527 nm $0.6180$ nm
Static bulk dielectric constant $\epsilon_{s,0}=\sqrt{\epsilon_{s,\parallel,0}\epsilon_{s,\bot,0}}$ [60] 12.0474 10.4743
Plasmon peak energy $\hbar\omega_{\text{pl}}$ [61] 22.0 eV 22.5 eV
Thomas-Fermi parameter $\alpha_{\text{TF}}$ (fit to ab-initio results [62]) 1.9 1.5
Interlayer gap $h$ [63, 64] 0.3 nm 0.3 nm
$1s$ spin-bright excitonic binding energy (calculated) 341.3-341.3 meV 361.2-361.2 meV
$1s$ spin-dark excitonic binding energy (calculated) 358.4-358.4 meV 369.2-369.2 meV
Radiative linewidth $\hbar\gamma^{b}_{\text{rad}}$ (calculated) 0.81 meV 0.71 meV
Coulomb enhancement $\sum_{\mathbf{q}}\varphi_{1s,\mathbf{q}}^{*\,K,\uparrow,\uparrow}$ (calculated) 0.9157 0.8677
Overlaps $\sum_{\mathbf{q}}\varphi_{1s,\mathbf{q}}^{*,K,\uparrow,\uparrow}\varphi_{1s,\mathbf{q}}^{K,\downarrow,\uparrow}$ (calculated) 0.9989 0.9998
Table 1: Parameters used in the numerical (Wannier equation) and analytical (optical spectra) calculations.
In Tab. 1, we display the parameters used in all analytical (optical spectra) and numerical (Wannier equation) calculations. In particular, we use the momentum-dependent model dielectric function from Ref. [65], which is fitted to ab-initio calculations from the Computational Materials Repository [62], and the screened Coulomb potential derived in Ref. [43] in the Wannier equation for the calculation of the excitonic binding energies and wave functions. To model the screening effect of the h-BN encapsulation, we fit the respective momentum-dependent dielectric function to ab-initio results [66] with static dielectric constant of bulk h-BN $\epsilon_{\text{h-BN},0}=\sqrt{\epsilon_{\text{h-BN},0,\parallel}\epsilon_{\text{h-BN},0,\bot}}=4.8$ [66]. For the light-matter interaction, we use the optical in-plane dielectric constant of h-BN $\epsilon_{\text{h-BN},\infty,\parallel}=4.87$ [67].